OpenStudy (anonymous):

pleaseeeeeeeeeeee helppppppppp Match the data set with the correct standard deviation. 80,81,84, 91, 99, 50, 93, 100 10,30,40, 80, 20, 50, 90, 110 100,310,400, 810, 290, 500, 930, 1100 Choose... 329.2 14.9 14.8 33.5

OpenStudy (anonymous):

are u there

OpenStudy (anonymous):

i will solve one set of data you can do the rest following the procedure 80,81,84, 91, 99, 50, 93, 100 1)first find the mean 2)then subtract each value from mean square it and add all of them 3) now divide by no of data points 4)take square root of the above and that's it you got SD! 1)find the mean \[\Large X=\frac{80+81+84+ 91+ 99+ 50 +93+ 100 }{8}=84.75\] subtract each value from it and square it\[(84.75-80)^2+(84.75-81)^2+(84.75-84)^2+(84.75-91)^2+(84.75-90)^2\] \[+(84.75-50)^2+(84.75-93)^2+(84.75-100)^2\] you get \[\Large 1787.5\] divide this by no of data points (n=8 in this case) \[\Large \frac{1787.5}{8}=223.5\] take square root and you will have SD \[\Large SD=\sqrt{223.43}=14.97\]

OpenStudy (anonymous):

you can find the others by similar way

OpenStudy (anonymous):

@brendas did you get it?

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

ok take the second data point and calculate its mean. let me know 10,30,40, 80, 20, 50, 90, 110

OpenStudy (anonymous):

53.75

OpenStudy (anonymous):

ok i need you to do now subtract each data point from it and square it then add all of them like i am doing first two terms \[\Large (53.75-10)^2+(53.75-30)^2+\] can you do this ? just subtract the values of data points from mean square them and add them let me know when you complete this .